OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 2x – 3y – z = –1 x – 4y + 3z = –11 2x + 9y + 3z = –13

OpenStudy (anonymous):

find x in terms of y and z from first equation 2x = -1 + 3y + z x = (-1 + 3y + z)/2

OpenStudy (anonymous):

Now put this value of x in second equation Would you like to do it @lornbeach

OpenStudy (whpalmer4):

Well, to isolate \(x\) in the second equation, I would add \(4y-3z\) to both sides, giving \[x=4y-3z-11\] Now substitute \(4y-3z-11\) wherever we encounter \(x\) in the other two equations: \[2(4y-3z-11) - 3y-z = -1\]\[2(4y-3z-11)+9y+3z=-13\] After distributing and collecting like terms, that would be: \[5y-7z=21\]\[17y-3z=9\]